OpenStudy (anonymous):

Compare each pair of numbers use <. =. > 3/4. 3/2

OpenStudy (anonymous):

3/4, 3/2 3/4=(3/4)(2/2)=6/8 3/2=(3/2)(4/4)=12/8 12/8 > 6/8 and 3/2 > 3/4 also 3/4 < 3/2 ans....